OpenStudy (anonymous):

1/2+3/4=? please help so i can understand

OpenStudy (anonymous):

5/4

OpenStudy (anonymous):

\[\frac{1}{2}+\frac{3}{4}=\]\[\frac{1\times2}{2\times2}+\frac{3}{4}=\]\[\frac{2}{4}+\frac{3}{4}=\]\[\frac{2+3}{4}=\]\[\frac{5}{4}\]